OpenStudy (anonymous):

Find the value of the polynomial -14-x^3-x^2 when x=-1 solve with steps please.

OpenStudy (anonymous):

-(-1)^3-(-1)^2-14=?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

what is -1^3?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

ok now -1*-1=1 correct

OpenStudy (anonymous):

so we have this correct? 1-(-1)^2-14=?

OpenStudy (anonymous):

it would be -13 right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

in my paper the options are -12,12,-16, or -14

OpenStudy (anonymous):

don't understand how it could be

OpenStudy (anonymous):

sorry -1^2=1 so 1-1=0 0-14=-14

OpenStudy (anonymous):

oh right ok that makes sense. thanks

OpenStudy (anonymous):

not a problem